Find the cofactor for given matrix "\begin{vmatrix}
 2& 5\\
 7 & -5
 
\end{vmatrix}"


1
Expert's answer
2021-11-03T17:36:50-0400

"M_{1,1}=(-1)^{1+1}\cdot (-5)=-5;M_{1,2}=(-1)^{1+2}\cdot 7=-7;\\
M_{2,1}=(-1)^{2+1}\cdot 5=-5; M_{2,2}=(-1)^{2+2}\cdot 2=2;\\"

Cofactor(A)="\begin{pmatrix}
 M_{1,1} & M_{1,2}\\
 M_{2,1} & M_{2,2}
\end{pmatrix}"Β ="\begin{pmatrix}
 -5 & -7\\
 -5 & 2
\end{pmatrix}"


More explanations:

"\begin{vmatrix}
 \colorbox{aqua}2 & \colorbox{aqua}5\\
 \colorbox{aqua} 7 & -5
\end{vmatrix}" "M_{1,1}=(-1)^{1+1}\cdot (-5)=-5"


"\begin{vmatrix}
 \colorbox{aqua}2 & \colorbox{aqua} 5\\
 7 &\colorbox{aqua} -5
\end{vmatrix}" "M_{1,2}=(-1)^{1+2}\cdot 7=-7"


"\begin{vmatrix}
 \colorbox{aqua}2 & 5\\
 \colorbox{aqua}7 &\colorbox{aqua} -5
\end{vmatrix}" "M_{2,1}=(-1)^{2+1}\cdot 5=-5"


"\begin{vmatrix}
2 & \colorbox{aqua}5\\
 \colorbox{aqua}7 &\colorbox{aqua} -5
\end{vmatrix}" "M_{2,2}=(-1)^{2+2}\cdot 2=2;"